# Mechanistic Indicators of Understanding in Large Language Models

### v4

@@ 2.2 The Linear Representation Hypothesis: Features as Directions
Given these reasons to think that LLMs learn features corresponding to concepts of entities and properties, the next question is how these features are encoded. A pivotal insight from MI is the linear representation hypothesis (LRH), which posits that features are encoded as directions within the model’s “latent space”—the highly multidimensional space (with as many dimensions as there are neurons in a layer) that is implicit or “latent” in each layer. The LRH is supported by many empirical (e.g., Mikolov et al., 2013; Radford et al., 2016) and theoretical arguments (see A.13 in Gurnee et al., 2023 as well as Bengio et al., 2013; Elhage et al., 2022) .
@@ 2.3 Superposition: Accommodating a World of Features
LLMs are trained on vast datasets and are expected to handle an enormous range of topics. This implies learning not just a few features, but a staggering number—features for countless entities (individual entities like “Mount Everest,” kinds like “mammal,” abstract entities like “justice”) and properties (attributes like “sharpness,” “redness,” or “sarcastic”). How can a model with a limited number of neurons in each layer accommodate this combinatorial explosion of features? The answer, MI research suggests, lies in superposition (Elhage et al., 2022). [^9] Superposition is an initial theory in the field of mechanistic interpretability that may yet be disproven. At the very least, however, it serves as a useful working hypothesis—much like caloric theory in its time (Henighan, 2024).
@@
This strategy comes with a trade-off, however: interference. If two features, A and B, are encoded in non-orthogonal directions, activating A also slightly activates B (see Fig. 2). [^10] This is the case of alternating interference. There is also simultaneous interference, where the joint activation of superposed features causes them to each activate more than if they were orthogonal (see A.6 in Gurnee et al., 2023).
@@
Two things explain why models can nevertheless compress many more features than one might first suspect. First, the number of near-orthogonal vectors that can be fitted in a space scales exponentially with its dimensionality (see the Johnson–Lindenstrauss lemma). With 100 dimensions, an LLM could “easily fit in 100,000 vectors at 88–92 degrees” (Sanderson, 2024), and current LLMs have vastly more dimensions than that.
@@
To “disentangle” these features, MI researchers have developed tools like sparse autoencoders (SAEs). [^11] Note that the reliability of sparse autoencoders is still debated in the literature. The central issue is whether SAEs actually discover a model’s features. In specific instances, SAEs exhibit limitations: features may fail to generalize beyond the training data (Sharkey et al., 2025), become misaligned with actual features through “absorption” or “splitting” (Chanin et al., 2024; Leask et al., 2025), suffer misinterpretation during manual inspection, or capture mere statistical artifacts (Heap et al., 2025). However, solutions for these issues are already being proposed (Makelov et al., 2024; Bussmann et al., 2025; Bayat et al., 2025). An SAE is an auxiliary neural network trained to decompress the dense, superimposed activation pattern of an LLM layer into a sparse, interpretable one where each active neuron corresponds to a single feature, rendering the activation pattern “monosemantic.” The SAE is trained to take the activation values of an LLM layer as input and to encode them in compressed form in a single hidden layer, but in such a way that the SAE can then still accurately decode its compressed representation to reproduce the original activation values (see Fig. 3).
@@
Researchers can then inspect these monosemantic neurons—for instance, by finding text inputs that maximally activate them—to understand what human-interpretable concepts the LLM has learned (Templeton et al., 2024). Thus, SAEs act like a computational “comb,” separating the entangled features in an LLM layer. It was using such an SAE that the “Golden Gate Bridge” feature and other features were discovered in state-of-the-art LLMs.
@@ 2.4 The Transformer Architecture: Deepening Conceptual Understanding
Having established that LLMs develop a rich repertoire of features (2.1), represent them as directions in latent space (2.2), and can store a vast number of them via superposition (2.3), the crucial next question is how these features are used to achieve relevant conceptual understanding in context. The transformer architecture plays a crucial role here by enabling the model to dynamically select and integrate information from different parts of an input sequence. Accordingly, we focus on decoder-only, causal masking, transformer-based LLMs—the standard architecture used notably in OpenAI’s GPT, Google’s Gemini, and Anthropic’s Claude models. [^12] This is documented for GPT-2/3 (Radford et al., 2019; Brown et al., 2020) and Gemini models (Gemini Team et al., 2025), and implied for Claude models in Anthropic’s research publications (e.g., for Haiku 3.5 Lindsey et al., 2025, , see Fig. 1). GPT-4/5 models have undisclosed architectures (OpenAI et al., 2024; OpenAI, 2025).
@@
In tracking a token’s journey through the model, the central pathway to focus on is the “residual stream” associated with that token (Elhage et al., 2021)—it constitutes the model’s continuously evolving understanding of a token from the initial input to the eventual output. This begins with a token entering the model and being converted into a numerical representation or vector—this is called embedding. This initial embedding is a first shot at interpreting the token, containing general information about it without considering the words around it. It amounts to giving the model a basic and entirely context-free definition that is useful for simple predictions, such as guessing what might follow the token based on common two-word patterns (“apple” is frequently followed by “pie”). Additionally, positional embeddings are added at this stage to encode information about the token’s position within the sequence.
@@
Each attention head learns to perform a specialized type of information retrieval from the residual streams of preceding tokens (see Fig. 5). [^13] Because transformer models are designed to process information in parallel to optimize training, however, an attention head at a certain depth within the network—say, at layer 3—can only draw on the residual streams of preceding tokens at the same depth within the network, i.e. at layer 3. This imposes counterintuitive constraints on the information flow: an attention head may sometimes be forced to compute information for itself that previous residual streams also compute, but only at later layers, thereby depriving the attention head of access to the relevant information (see Nostalgebraist, 2024a). The exchange of information does not happen at random, however. Through the attention mechanism, the head determines what information to seek and identifies which residual streams contain it. The attention head reads from residual streams in three distinct ways. First, from the current stream, it forms a query specifying the kind of information it is looking for. Second, it computes “keys” for all token positions in the context, which advertise the kind of information available at those positions. When a query from the current token strongly matches a key from a previous token, a high “attention score” is generated. [^14] This attention score is obtained by first computing the dot product between the query and key vectors, followed by a softmax (see Fig. 5). These two operations introduce crucial non-linearities into the attention mechanism—enabling, as in the activation functions used in MLPs, a form of selectivity. Third, this score then determines how much of the “value”—the actual information available to be passed along—from that previous token’s residual stream is copied and added to the current token’s stream.
@@
On the other hand, MLPs augment the resulting combination with associated knowledge recalled from the model’s learned parameters (Nanda et al., 2023c; Geva et al., 2023; Chughtai et al., 2024). This information recall can take many forms. But the one that is most fundamental to conceptual understanding is the ability to connect a strongly activated feature to other features that help define and identify it. Thus, if the “cat” feature is strongly activated within the current residual stream, an MLP layer might recall associated information stored within its weights, enabling the LLM to connect the “cat” feature to other features that are definitive of being a cat, such as “mammal,” “pointy-earedness,” and “furriness” (we explain how such feature-activated information recall works in section 3.1).
This recall process significantly enriches the model’s features, making them more nuanced and useful for predicting subsequent tokens that depend on the ability to exploit these conceptual connections. This is fundamental to an LLM’s ability to move beyond superficial token-co-occurrence patterns and leverage conceptual connections among features. And grasping these conceptual connections amounts to a form of understanding. “To have understanding,” Jonathan Kvanvig writes, is notably “to grasp … conceptual connections” (2018, p. 699). There is a rich philosophical tradition holding that to understand a concept is, above all, to master its connections to other concepts that obtain in virtue of the content of the concept—one does not grasp the concept “bachelor” unless one grasps its connections to the concepts “man” and “unmarried.” [^18] Such conceptual connections figure particularly prominently in the inferentialist tradition (Sellars, Dummett, Harman, Brandom, Block, and others, which has been thought to be especially well-suited to making sense of LLMs (Piantadosi and Hill, 2022; Piantadosi et al., 2024). But conceptual connections are also key to the much older idea of “grammatical truths” (Wittgenstein), “analytic truths” (Kant), or “truths of reason” (Leibniz). By empowering LLMs to recall analogous connections between features, MLPs significantly strengthen the case for the attribution of something like conceptual understanding to LLMs. Moreover, both the combination of features and the recall of associated information often occur in superposition within the MLP layers, further leveraging the model’s dense representational capacity.
@@
This contrasts with classical multi-layer perceptrons without skip connections, where activations are thoroughly and irretrievably transformed from one layer to the next. This might lead to a loss of important original information, which makes it harder for the network to learn (this is related to the “vanishing gradient problem”). Preserving the original information that enters a transformer block avoids this problem, giving the network the freedom to pass information through a transformer block unchanged where necessary. This gives the model the freedom to be selective about which layers it processes information through instead of having every token be affected by every layer. And in fact, it has been observed that MLPs and attention layers are highly selective, each interacting with a narrow, dedicated subspace of the residual stream—introducing a kind of division of labor (Elhage et al., 2021; Su et al., 2025). The model can pick up on more complex patterns as a result.
@@
The assumption of a single, fixed latent space is supported by an MI technique called “logit lens” (Nostalgebraist, 2024b), which works by treating an activation vector in an intermediate layer as if it were the final one and running it through the unembedding matrix to see what next-token predictions would be made based on the information accumulated up to that point. By applying the logit lens at each layer, researchers can create snapshots of the model’s evolving predictions. They found that the model did not produce radically different predictions at each layer, but successively refined them, supporting the view of the residual stream as a cumulative computational workspace.
@@ 3  State-of-the-World Understanding
State-of-the-world understanding involves the LLM forming an internal model of how distinct features relate to each other to constitute facts. This allows the model to learn factual associations between features and output sentences like “Marie Curie was a physicist,” not merely as statistically likely sequences of tokens, but as outward expressions of an internal organization that links these features. This ability to learn factual associations between features aligns with a central tenet in the philosophical literature on understanding: that understanding is factive, as it relies on connections that obtain in the actual world (Baumberger et al., 2017).
@@ 3.1 Learning Factual Connections
This highlights the capacity of LLMs to recall associated information via factual as well as conceptual associations. Traditionally, deep learning models have been seen as cascades of feature combinations, successively assembling lower-level features into higher-level ones. The Convolutional Neural Networks used for image recognition provide the paradigmatic example of this—first detecting lines, edges and simple patterns, then assembling them into more complex patterns, and eventually arriving at objects. In LLMs, MLPs do act as such feature-combiners (e.g. when combining the “Michael” and “Jordan” features); but they also go significantly beyond this, serving as feature-activated information recallers that respond to the strong activation of a feature by augmenting and enriching it with the conceptual and factual information stored in their parameters.
@@
This mechanism of factual recall through MLP layers has been empirically validated by MI research. Studies using targeted interventions confirm that MLP layers are the primary locus of information enrichment (Geva et al., 2023; Nanda et al., 2023c). There is also strong evidence suggesting that information typically gets recalled in the residual stream of the subject—in our example, the “Golden Gate Bridge” stream—thereby making it available for upcoming residual streams that might need it (Geva et al., 2023; Nanda et al., 2023c; Ameisen et al., 2025; Lindsey et al., 2025). Furthermore, as we mentioned in footnote 10, attention heads also recall information during their routing task, via their direct linear projections: the value projection and the final output projection (Geva et al., 2023; Chughtai et al., 2024).
@@
MLPs might also handle predication involving temporal properties using conditional gating of recall: the model can learn to retrieve a fact only when certain conditions are met. For example, the activation of a “year 1900” feature could inhibit the usual association of “Golden Gate Bridge” with “San Francisco,” allowing the model to express that the bridge was not in San Francisco in 1900. To our knowledge, this specific temporal-gating hypothesis has yet to be tested, but there is evidence that LLMs encode time-related features (Gurnee and Tegmark, 2024) and often implement more complex cases of predication via feature-level inhibition (e.g., Lindsey et al., 2025, §8)
@@ 3.2 Internal Models of External States: The Othello-GPT Findings
Research on Othello-GPT provides a compelling affirmative example. Othello-GPT is a language model based on GPT-2 that was trained to generate legal moves in the game of Othello by predicting the next move. The model was trained solely on game transcriptions, without ever seeing the board or even the game’s rules. Despite this limited input, when provided with a sequence of moves from a game, Othello-GPT reliably outputs only moves that are legal for the current board position. This ability generalizes to unseen sequences. The model was tested on game sequences that had been systematically withheld from the training data (e.g., all the sequences downstream of the opening move C5, which represents a quarter of the game tree). Even when traversing these unfamiliar branches of the game tree, the model successfully predicted legal moves, demonstrating that it was not simply relying on memorized game sequences (Li et al., 2023).
@@
To determine whether Othello-GPT was actually maintaining an internal representation of the current state of the board—an “emergent world model,” as Li et al. (2023) call it—the researchers employed a technique known as probing. Probing, established as a key method for mechanistic interpretability by Alain and Bengio (2017), serves to detect if and where a given feature is encoded in a model. [^20] For a philosophical discussion of when linear probes reveal LLMs’ beliefs, see Herrmann and Levinstein (2024). One starts by selecting a target feature and a layer to investigate. One then trains a separate model (the “probe”) to predict, based on the chosen layer’s activation pattern when presented with an input, whether the feature is present in that input. The probe learns to do this by seeing lots of examples of both activation patterns produced by inputs explicitly labeled as containing the feature and activation patterns produced by inputs explicitly labeled as lacking the feature. As readers of Wittgenstein (1953, §§28–33) on the indeterminacy of ostensive definitions will know, there is a challenge involved in selecting positive and negative examples that unambiguously exemplify the target feature as opposed to its many correlates: a collection of inputs mentioning politicians instantiate the feature “politician,” but the contrast class of negative examples needs to be carefully chosen to ensure that the probe does not learn to probe for the “political” feature, or the “person” feature, which are equally instantiated by the positive examples (see Gurnee et al., 2023). Once suitable training data have been assembled and labeled, however, one can pass the inputs through the model to be investigated, gather the activation patterns at the chosen layer, and see if the probe can learn to predict the feature’s presence from these activation patterns alone. If the probe can learn to do this accurately on new, unseen inputs, this provides strong evidence that the feature is encoded in that layer.
@@
Li et al. trained probes to predict the state of each of the 64 board squares—whether it was empty, black, or white—solely from the model’s internal activations. Their initial attempts using linear probes yielded high error rates. This ran counter to the Linear Representation Hypothesis (LRH), suggesting that the model did not store the board state in a linear, easily intelligible format. However, when they switched to non-linear probes (specifically, 2-layer multilayer perceptrons), accuracy soared, with error rates dropping as low as $1.7\%$. This indicated that a representation of the board was indeed encoded within the network, albeit in a complex form.
Yet the initial failure to find simple linear encodings of the board state left the door open to a more skeptical interpretation: if a model’s internal states are highly entangled and non-linear, a skeptic might argue that it has not really learned to represent the board. It may merely have found a way to compress the statistical regularities of the move sequences. On this view, the model might simply be factorizing the probability distribution of the data using uninterpretable latent variables—internal states that happen to correlate with board features because they are instrumentally useful for prediction, but which do not share the logical structure of the game. Just as a weather model might have variables that correlate with “jet streams” without understanding meteorology, an LLM might track obscure statistical dependencies that are output-equivalent to understanding the board state, yet remain a tangled and fundamentally unintelligible compression of the data, more akin to a collection of heuristics than to emergent state-of-the-world understanding. [^21] We are indebted to an anonymous reviewer for this way of phrasing the skeptical challenge.
@@
This skeptical challenge became increasingly unconvincing in light of subsequent analyses, however. A study by Nanda et al. (2023b) provided the first counter-argument, revealing that the model’s internal representation of the board was simpler and more elegant than first thought. The trick was to reframe what to probe for. Instead of looking for features describing the board in absolute terms (“Black” vs. “White” discs), Nanda et al. hypothesized that the model might track pieces in relation to a player’s perspective (“Mine” vs. “Yours”). This time, linear probes were successful. To appreciate the import of this result for the skeptical challenge, one needs to understand how linear probes work.
@@
Unlike the non-linear probes used by Li et al., linear probes operate under the assumption that the LRH holds true: whatever features the model learns are represented as lines or directions in its latent space. To visualize this, imagine the entire activation pattern that an input produces at one layer not as a list of neuron activations or numbers, but as a single point: the point picked out by the “activation vector” that this list of numbers defines. If a model has learned to treat a direction as corresponding to a feature, it will organize the inputs so that inputs with that feature register as a point in one region of latent space and inputs without that feature register as a point in a clearly distinct region.

### v5

@@ 2.2 The Linear Representation Hypothesis: Features as Directions
Given these reasons to think that LLMs learn features corresponding to concepts of entities and properties, the next question is how these features are encoded. A pivotal insight from MI is the linear representation hypothesis (LRH), which posits that features are encoded as directions within the model’s “latent space”—the highly multidimensional space (with as many dimensions as there are neurons in a layer) that is implicit or “latent” in each layer. The LRH is supported by many empirical (e.g., Mikolov et al.,, 2013; Radford et al.,, 2016) and theoretical arguments (see A.13 in Gurnee et al.,, 2023 as well as Bengio et al.,, 2013; Elhage et al.,, 2022) .
@@ 2.3 Superposition: Accommodating a World of Features
LLMs are trained on vast datasets and are expected to handle an enormous range of topics. This implies learning not just a few features, but a staggering number—features for countless entities (individual entities like “Mount Everest,” kinds like “mammal,” abstract entities like “justice”) and properties (attributes like “sharpness,” “redness,” or “sarcastic”). How can a model with a limited number of neurons in each layer accommodate this combinatorial explosion of features? The answer, MI research suggests, lies in superposition (Elhage et al.,, 2022). [^9] Superposition is an initial theory in the field of mechanistic interpretability that may yet be disproven. At the very least, however, it serves as a useful working hypothesis—much like caloric theory in its time (Henighan,, 2024).
@@
This strategy comes with a trade-off, however: interference. If two features, A and B, are encoded in non-orthogonal directions, activating A also slightly activates B (see Fig. 2). [^10] This is the case of alternating interference. There is also simultaneous interference, where the joint activation of superposed features causes them to each activate more than if they were orthogonal (see A.6 in Gurnee et al.,, 2023).
@@
Two things explain why models can nevertheless compress many more features than one might first suspect. First, the number of near-orthogonal vectors that can be fitted in a space scales exponentially with its dimensionality (see the Johnson–Lindenstrauss lemma). With 100 dimensions, an LLM could “easily fit in 100,000 vectors at 88–92 degrees” (Sanderson,, 2024), and current LLMs have vastly more dimensions than that.
@@
To “disentangle” these features, MI researchers have developed tools like sparse autoencoders (SAEs). [^11] Note that the reliability of sparse autoencoders is still debated in the literature. The central issue is whether SAEs actually discover a model’s features. In specific instances, SAEs exhibit limitations: features may fail to generalize beyond the training data (Sharkey et al.,, 2025), become misaligned with actual features through “absorption” or “splitting” (Chanin et al.,, 2024; Leask et al.,, 2025), suffer misinterpretation during manual inspection, or capture mere statistical artifacts (Heap et al.,, 2025). However, solutions for these issues are already being proposed (Makelov et al.,, 2024; Bussmann et al.,, 2025; Bayat et al.,, 2025). An SAE is an auxiliary neural network trained to decompress the dense, superimposed activation pattern of an LLM layer into a sparse, interpretable one where each active neuron corresponds to a single feature, rendering the activation pattern “monosemantic.” The SAE is trained to take the activation values of an LLM layer as input and to encode them in compressed form in a single hidden layer, but in such a way that the SAE can then still accurately decode its compressed representation to reproduce the original activation values (see Fig. 3).
@@
Researchers can then inspect these monosemantic neurons—for instance, by finding text inputs that maximally activate them—to understand what human-interpretable concepts the LLM has learned (Templeton et al.,, 2024). Thus, SAEs act like a computational “comb,” separating the entangled features in an LLM layer. It was using such an SAE that the “Golden Gate Bridge” feature and other features were discovered in state-of-the-art LLMs.
@@ 2.4 The Transformer Architecture: Deepening Conceptual Understanding
Having established that LLMs develop a rich repertoire of features (2.1), represent them as directions in latent space (2.2), and can store a vast number of them via superposition (2.3), the crucial next question is how these features are used to achieve relevant conceptual understanding in context. The transformer architecture plays a crucial role here by enabling the model to dynamically select and integrate information from different parts of an input sequence. Accordingly, we focus on decoder-only, causal masking, transformer-based LLMs—the standard architecture used notably in OpenAI’s GPT, Google’s Gemini, and Anthropic’s Claude models. [^12] This is documented for GPT-2/3 (Radford et al.,, 2019; Brown et al.,, 2020) and Gemini models (Gemini Team et al.,, 2025), and implied for Claude models in Anthropic’s research publications (e.g., for Haiku 3.5 Lindsey et al.,, 2025, , see Fig. 1). GPT-4/5 models have undisclosed architectures (OpenAI et al.,, 2024; OpenAI,, 2025).
@@
In tracking a token’s journey through the model, the central pathway to focus on is the “residual stream” associated with that token (Elhage et al.,, 2021)—it constitutes the model’s continuously evolving understanding of a token from the initial input to the eventual output. This begins with a token entering the model and being converted into a numerical representation or vector—this is called embedding. This initial embedding is a first shot at interpreting the token, containing general information about it without considering the words around it. It amounts to giving the model a basic and entirely context-free definition that is useful for simple predictions, such as guessing what might follow the token based on common two-word patterns (“apple” is frequently followed by “pie”). Additionally, positional embeddings are added at this stage to encode information about the token’s position within the sequence.
@@
Each attention head learns to perform a specialized type of information retrieval from the residual streams of preceding tokens (see Fig. 5). [^13] Because transformer models are designed to process information in parallel to optimize training, however, an attention head at a certain depth within the network—say, at layer 3—can only draw on the residual streams of preceding tokens at the same depth within the network, i.e. at layer 3. This imposes counterintuitive constraints on the information flow: an attention head may sometimes be forced to compute information for itself that previous residual streams also compute, but only at later layers, thereby depriving the attention head of access to the relevant information (see Nostalgebraist, 2024a, ). The exchange of information does not happen at random, however. Through the attention mechanism, the head determines what information to seek and identifies which residual streams contain it. The attention head reads from residual streams in three distinct ways. First, from the current stream, it forms a query specifying the kind of information it is looking for. Second, it computes “keys” for all token positions in the context, which advertise the kind of information available at those positions. When a query from the current token strongly matches a key from a previous token, a high “attention score” is generated. [^14] This attention score is obtained by first computing the dot product between the query and key vectors, followed by a softmax (see Fig. 5). These two operations introduce crucial non-linearities into the attention mechanism—enabling, as in the activation functions used in MLPs, a form of selectivity. Third, this score then determines how much of the “value”—the actual information available to be passed along—from that previous token’s residual stream is copied and added to the current token’s stream.
@@
On the other hand, MLPs augment the resulting combination with associated knowledge recalled from the model’s learned parameters (Nanda et al., 2023c, ; Geva et al.,, 2023; Chughtai et al.,, 2024). This information recall can take many forms. But the one that is most fundamental to conceptual understanding is the ability to connect a strongly activated feature to other features that help define and identify it. Thus, if the “cat” feature is strongly activated within the current residual stream, an MLP layer might recall associated information stored within its weights, enabling the LLM to connect the “cat” feature to other features that are definitive of being a cat, such as “mammal,” “pointy-earedness,” and “furriness” (we explain how such feature-activated information recall works in section 3.1).
This recall process significantly enriches the model’s features, making them more nuanced and useful for predicting subsequent tokens that depend on the ability to exploit these conceptual connections. This is fundamental to an LLM’s ability to move beyond superficial token-co-occurrence patterns and leverage conceptual connections among features. And grasping these conceptual connections amounts to a form of understanding. “To have understanding,” Jonathan Kvanvig writes, is notably “to grasp … conceptual connections” (2018, p. 699). There is a rich philosophical tradition holding that to understand a concept is, above all, to master its connections to other concepts that obtain in virtue of the content of the concept—one does not grasp the concept “bachelor” unless one grasps its connections to the concepts “man” and “unmarried.” [^18] Such conceptual connections figure particularly prominently in the inferentialist tradition (Sellars, Dummett, Harman, Brandom, Block, and others, which has been thought to be especially well-suited to making sense of LLMs (Piantadosi and Hill,, 2022; Piantadosi et al.,, 2024). But conceptual connections are also key to the much older idea of “grammatical truths” (Wittgenstein), “analytic truths” (Kant), or “truths of reason” (Leibniz). By empowering LLMs to recall analogous connections between features, MLPs significantly strengthen the case for the attribution of something like conceptual understanding to LLMs. Moreover, both the combination of features and the recall of associated information often occur in superposition within the MLP layers, further leveraging the model’s dense representational capacity.
@@
This contrasts with classical multi-layer perceptrons without skip connections, where activations are thoroughly and irretrievably transformed from one layer to the next. This might lead to a loss of important original information, which makes it harder for the network to learn (this is related to the “vanishing gradient problem”). Preserving the original information that enters a transformer block avoids this problem, giving the network the freedom to pass information through a transformer block unchanged where necessary. This gives the model the freedom to be selective about which layers it processes information through instead of having every token be affected by every layer. And in fact, it has been observed that MLPs and attention layers are highly selective, each interacting with a narrow, dedicated subspace of the residual stream—introducing a kind of division of labor (Elhage et al.,, 2021; Su et al.,, 2025). The model can pick up on more complex patterns as a result.
@@
The assumption of a single, fixed latent space is supported by an MI technique called “logit lens” (Nostalgebraist, 2024b, ), which works by treating an activation vector in an intermediate layer as if it were the final one and running it through the unembedding matrix to see what next-token predictions would be made based on the information accumulated up to that point. By applying the logit lens at each layer, researchers can create snapshots of the model’s evolving predictions. They found that the model did not produce radically different predictions at each layer, but successively refined them, supporting the view of the residual stream as a cumulative computational workspace.
@@ 3  State-of-the-World Understanding
State-of-the-world understanding involves the LLM forming an internal model of how distinct features relate to each other to constitute facts. This allows the model to learn factual associations between features and output sentences like “Marie Curie was a physicist,” not merely as statistically likely sequences of tokens, but as outward expressions of an internal organization that links these features. This ability to learn factual associations between features aligns with a central tenet in the philosophical literature on understanding: that understanding is factive, as it relies on connections that obtain in the actual world (Baumberger et al.,, 2017).
@@ 3.1 Learning Factual Connections
This highlights the capacity of LLMs to recall associated information via factual as well as conceptual associations. Traditionally, deep learning models have been seen as cascades of feature combinations, successively assembling lower-level features into higher-level ones. The Convolutional Neural Networks used for image recognition provide the paradigmatic example of this—first detecting lines, edges and simple patterns, then assembling them into more complex patterns, and eventually arriving at objects. In LLMs, MLPs do act as such feature-combiners (e.g. when combining the “Michael” and “Jordan” features); but they also go significantly beyond this, serving as feature-activated information recallers that respond to the strong activation of a feature by augmenting and enriching it with the conceptual and factual information stored in their parameters (Beckmann,, 2026).
@@
This mechanism of factual recall through MLP layers has been empirically validated by MI research. Studies using targeted interventions confirm that MLP layers are the primary locus of information enrichment (Geva et al.,, 2023; Nanda et al., 2023c, ). There is also strong evidence suggesting that information typically gets recalled in the residual stream of the subject—in our example, the “Golden Gate Bridge” stream—thereby making it available for upcoming residual streams that might need it (Geva et al.,, 2023; Nanda et al., 2023c, ; Ameisen et al.,, 2025; Lindsey et al.,, 2025). Furthermore, as we mentioned in footnote 17, attention heads also recall information during their routing task, via their direct linear projections: the value projection and the final output projection (Geva et al.,, 2023; Chughtai et al.,, 2024).
@@
MLPs might also handle predication involving temporal properties using conditional gating of recall: the model can learn to retrieve a fact only when certain conditions are met. For example, the activation of a “year 1900” feature could inhibit the usual association of “Golden Gate Bridge” with “San Francisco,” allowing the model to express that the bridge was not in San Francisco in 1900. To our knowledge, this specific temporal-gating hypothesis has yet to be tested, but there is evidence that LLMs encode time-related features (Gurnee and Tegmark,, 2024) and often implement more complex cases of predication via feature-level inhibition (e.g., Lindsey et al., 2025, §8)
@@ 3.2 Internal Models of External States: The Othello-GPT Findings
Research on Othello-GPT provides a compelling affirmative example. Othello-GPT is a language model based on GPT-2 that was trained to generate legal moves in the game of Othello by predicting the next move. The model was trained solely on game transcriptions, without ever seeing the board or even the game’s rules. Despite this limited input, when provided with a sequence of moves from a game, Othello-GPT reliably outputs only moves that are legal for the current board position. This ability generalizes to unseen sequences. The model was tested on game sequences that had been systematically withheld from the training data (e.g., all the sequences downstream of the opening move C5, which represents a quarter of the game tree). Even when traversing these unfamiliar branches of the game tree, the model successfully predicted legal moves, demonstrating that it was not simply relying on memorized game sequences (Li et al.,, 2023).
@@
To determine whether Othello-GPT was actually maintaining an internal representation of the current state of the board—an “emergent world model,” as Li et al. (2023) call it—the researchers employed a technique known as probing. Probing, established as a key method for mechanistic interpretability by Alain and Bengio (2017), serves to detect if and where a given feature is encoded in a model. [^20] For a philosophical discussion of when linear probes reveal LLMs’ beliefs, see Herrmann and Levinstein (2024). One starts by selecting a target feature and a layer to investigate. One then trains a separate model (the “probe”) to predict, based on the chosen layer’s activation pattern when presented with an input, whether the feature is present in that input. The probe learns to do this by seeing lots of examples of both activation patterns produced by inputs explicitly labeled as containing the feature and activation patterns produced by inputs explicitly labeled as lacking the feature. As readers of Wittgenstein (1953, §§28–33) on the indeterminacy of ostensive definitions will know, there is a challenge involved in selecting positive and negative examples that unambiguously exemplify the target feature as opposed to its many correlates: a collection of inputs mentioning politicians instantiate the feature “politician,” but the contrast class of negative examples needs to be carefully chosen to ensure that the probe does not learn to probe for the “political” feature, or the “person” feature, which are equally instantiated by the positive examples (see Gurnee et al.,, 2023). Once suitable training data have been assembled and labeled, however, one can pass the inputs through the model to be investigated, gather the activation patterns at the chosen layer, and see if the probe can learn to predict the feature’s presence from these activation patterns alone. If the probe can learn to do this accurately on new, unseen inputs, this provides strong evidence that the feature is encoded in that layer.
@@
Li et al., trained probes to predict the state of each of the 64 board squares—whether it was empty, black, or white—solely from the model’s internal activations. Their initial attempts using linear probes yielded high error rates. This ran counter to the Linear Representation Hypothesis (LRH), suggesting that the model did not store the board state in a linear, easily intelligible format. However, when they switched to non-linear probes (specifically, 2-layer multilayer perceptrons), accuracy soared, with error rates dropping as low as $1.7\%$. This indicated that a representation of the board was indeed encoded within the network, albeit in a complex form.
Yet the initial failure to find simple linear encodings of the board state left the door open to a more skeptical interpretation: if a model’s internal states are highly entangled and non-linear, a skeptic might argue that it has not really learned to represent the board at all. Consider an analogy with jet streams in a climate model. Even if no explicitly represented variables stand in for jet streams, they are still present in an emergent way in the dynamics—recoverable, in principle, by a classifier informed by the relevant background knowledge. Yet one would not want to say that the climate model “understands” jet streams. In the same way, a deep neural network may lack any explicit representation of the board state, even if that state can be decoded from its distributed activation patterns. The skeptic’s point is that if one would not say that the climate model “understands” jet streams, one should not say that LLMs “understand” the board state in Othello either. [^21] We are indebted to an anonymous reviewer for this way of phrasing the skeptical challenge.
@@
This skeptical challenge became increasingly unconvincing in light of subsequent analyses, however. A study by Nanda et al. () provided the first counter-argument, revealing that the model’s internal representation of the board was simpler and more elegant than first thought. The trick was to reframe what to probe for. Instead of looking for features describing the board in absolute terms (“Black” vs. “White” discs), Nanda et al. hypothesized that the model might track pieces in relation to a player’s perspective (“Mine” vs. “Yours”). This time, linear probes were successful. To appreciate the import of this result for the skeptical challenge, one needs to understand how linear probes work.
@@
Unlike the non-linear probes used by Li et al.,, linear probes operate under the assumption that the LRH holds true: whatever features the model learns are represented as lines or directions in its latent space. To visualize this, imagine the entire activation pattern that an input produces at one layer not as a list of neuron activations or numbers, but as a single point: the point picked out by the “activation vector” that this list of numbers defines. If a model has learned to treat a direction as corresponding to a feature, it will organize the inputs so that inputs with that feature register as a point in one region of latent space and inputs without that feature register as a point in a clearly distinct region.